# Measurement as Bricolage: Examining How Data Scientists Construct Target Variables for Predictive Modeling Tasks

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proc. ACM Hum.-Comput. Interact. 5, CSCW2; 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learningccs: Computing methodologies Supervised learningjournal: PACMHCIjournalvolume: 7journalnumber: CSCW2article: 1publicationmonth: 11journalyear: 2025doi: 10.1145/XXXXXXX
@@ 1. Introduction
Figure 1. An illustration of the relationship between the target variable and problem formulation. A prediction task describes a mapping from features to a target variable. A problem formulation connects a prediction task to high-level modeling goals, including factors such as how risk scores are converted to intervention recommendations and the target population at deployment time.
One critical aspect of problem formulation is the process of target variable construction (Figure 1). Many outcomes of interest to data scientists — such as the “authenticity” of student writing or the “healthcare need” of patients — are latent, theoretical constructs that are not directly observable in data. To develop a model targeting these constructs, data scientists must construct a proxy variable based on outcomes that are readily available. For example, a data scientist might choose to predict patients’ cost of medical care or their number of recently diagnosed chronic health conditions as a proxy for their “healthcare need” (Obermeyer et al., 2019). Traditionally, measurement theory assumes that analysts follow a top-down workflow while mapping unobserved constructs to measurements in data (Jacobs and Wallach, 2021). For example, a learning scientist might identify a latent construct they would like to measure among students (e.g., “math proficiency”) and design an operationalization (e.g., a test) before collecting data (e.g., student test scores). Disciplines in the quantitative social sciences have developed a rich set of guidelines (e.g., (Van’t Veer and Giner-Sorolla, 2016; Miller and Salkind, 2002)) and modeling tools (e.g., (Everett, 2013; Singh, 2007)) to help analysts implement this top-down measurement workflow.
@@
To understand data scientists’ current practices and challenges in constructing target variables—and to identify opportunities to better support target variable construction in practice—in this work, we conducted semi-structured interviews with fifteen data scientists from education (N=8) and healthcare (N=7). We asked data scientists to walk us through how they navigate target variable construction in prior modeling projects. Participants also engaged in a target variable construction task, allowing us to observe their reasoning in a new modeling context. Our goal in these interviews was to understand how data scientists intertwine two conceptually distinct activities, measurement of latent constructs and prediction of target variables, during problem formulation (Mussgnug, 2022). While prior empirical studies of data science practice have often centered prediction (e.g., (Alspaugh et al., 2018; Wongsuphasawat et al., 2019; Kandel et al., 2012)), we broaden our scope of inquiry to investigate the interplay between measurement validity and predictive utility in problem formulation.
@@
Based on our findings, we synthesize a new process model that better reflects data scientists’ bricolage approach towards measurement (Figure 2). Data scientists balance the validity of their formulation with other criteria such as predictive performance, simplicity, portability, and resource requirements. Data scientists narrow in on a suitable formulation by drawing upon a rich repertoire of problem (re)formulation strategies. Data scientists apply a strategy, take a step back to assess their progress along each criterion, and then apply a new strategy until all evaluation criteria are satisfied. Data scientists sometimes discontinue their projects if they conclude that no strategies are sufficient given the resources at hand.
@@
  • Empirical findings from interviews with fifteen data scientists in healthcare and education, synthesized into a process model that reveals how practitioners weigh competing criteria ($\S$§ 4.1), apply five strategic approaches to (re)formulation ($\S$§ 4.2), and evaluate validity through both theory-driven and data-driven practices ($\S$§ 4.3).
  • Design implications for how the CSCW, HCI, and ML research communities might better scaffold data scientists’ bricolage practices going forward. These recommendations explore how we might enhance the validity of data scientists’ resulting formulations while also supporting the flexibility that makes bricolage effective ($\S$§ 5).
@@ 2. Background and Related Work
In this section, we begin by introducing Claude Lévi-Strauss’ theory of Bricolage. We then describe existing Science and Technology Studies (STS), HCI, and CSCW research that has investigated questions related to categorization, measurement, and target variable construction in data science practice. We conclude by providing an overview of key concepts from measurement theory in the quantitative social sciences which we build upon in this work.
@@ 2.1. Bricolage and the Science of the Concrete
Bricolage, defined by Claude Lévi-Strauss as “making the most of available resources”(Levi-Strauss, 1966), describes how actors (i.e., bricoleurs) construct artifacts in creative and adaptive ways. Bricolage is central to what Lévi-Strauss calls the science of the concrete, which contrasts with modern scientific thought by focusing on the materials directly at hand rather than abstract principles. Bricolage contains three elements: repertoire, dialogue, outcome (Duymedjian and Rüling, 2010). Repertoire describes the resources—physical artifacts, ideas, or skills—that the bricoleur brings to bear on a problem. Whereas Lévi-Strauss viewed the traditional engineer or scientist as acquiring resources tailor-made to specification, the bricoleur slowly assembles various “odds and ends” that might be useful in the future. Then, when faced with a new problem, the bricoleur engages in dialogue with these materials—a process of arranging and re-arranging materials in her repertoire to find a workable solution. While engaging in dialogue, the bricoleur adopts old objects for new purposes and combines concepts with materials in novel ways. Finally, outcome describes the artifact resulting from the bricolage process. While bricolage entails a tolerance for imperfection, Lévi-Strauss argues that bricoleurs sometimes identify “brilliant unforeseen results” (Levi-Strauss, 1966) as part of their process.
Prior research has used Lévi-Strauss’ framework to describe many forms of knowledge work as bricolage (e.g., entrepreneurship (Mateus and Sarkar, 2024; Duymedjian and Rüling, 2010), design (Büscher et al., 2001; Louridas, 1999; Vallgårda and Fernaeus, 2015), management (Duymedjian and Rüling, 2010), and programming (Turkle, 2011)). For instance, Baker and Nelson (2005) use bricolage to explain how entrepreneurs build business ventures under resource constraints. They observe that entrepreneurial success hinges on the ability of firms to combine their limited resources in creative, adaptive ways.
@@
Turkle describes this style of tinkering as important for developing an intuition for the complexity of computer programs. Like a painter who steps back to inspect a canvas between brush strokes, programmers try one approach, re-evaluate, and then try another. Turkle observes that this process often leads programmers to more elegant solutions than could be imagined through the more traditional process of structured planning.
In this work, we draw a novel connection between data science work and the practice of bricolage. Much like an entrepreneur, programmer, or painter, data scientists construct target variables by combining existing resources in creative, adaptive ways. Instead of instantiating abstract plans through top-down design—akin to the structured programming approach taught in classrooms—data scientists engage in more direct dialogue with their data and modeling tools, making do with the materials at hand. And analogous to past research investigating how to support such exploratory and improvisational coding practices (Kery and Myers, 2017), this has implications for the design of practical tools to support target variable construction in practice. We highlight connections between target variable construction and bricolage throughout the findings ($\S$§ 4) and describe implications for tool design in the discussion ($\S$§ 5).
@@ 2.2. Tracing the Social Construction of Target Variables
Target variable construction also relates to a rich body of literature in Science and Technology Studies (STS) and philosophy of science that investigates the social processes that give rise to categories in data (Latour and Woolgar, 2013; Bowker, 2000; Pine and Liboiron, 2015; Harvey et al., 2024; Porter, 1996; Crosby, 1997; Hacking, 1999; Mol, 2002; Suchman, 1993). Latour and Woolgar (2013) provide an account of the social construction of scientific facts within laboratories, highlighting that facts are not merely discovered but actively produced through a complex interplay of social and technical factors. Likewise, Bowker (2000) examines how systems of classification, such as the International Classification of Diseases (ICD), shape and are shaped by the social and political contexts in which they are embedded. Building on Bowker’s work, recent studies have also explored the construction of racial and ethnic categories within AI models (e.g., (Abdu et al., 2023; Mickel, 2024)) and charted measurement assumptions underpinning motion capture technologies (Harvey et al., 2024).
Most related to our study, Muller et al. (2021) interviewed fifteen data scientists to investigate the data labeling and annotation practices of data science teams. Muller et al. (2021) identified three approaches that teams used while defining labels: principled, iterative, and improvisational. Teams engaging in principled design followed a top-down planning approach, systematically specifying annotation guidelines before data labeling. Teams leveraging iterative design took a bottom-up approach, incorporating opportunities for annotation protocol refinement throughout the labeling process. Finally, teams engaging in improvisational design viewed data labeling as an open-ended, exploitative process. Our study investigates a distinct question of how data science teams construct target variables for predictive modeling tasks. Rather than studying how data science teams design “ground truth” labels anew, we examine how data scientists connect their existing data to their high-level modeling objectives as part of the problem formulation process.
@@ 2.3. Empirical Studies of Data Science Practice
However, we currently have an incomplete understanding of how data scientists construct target variables while formulating prediction tasks. To date, Passi and Barocas (2019) provide the most detailed account of how problem formulation unfolds in real-world modeling contexts. Over the course of an eighteen-month ethnographic study of a corporate data science team, Passi and Barocas (2019) observed that data scientists repeatedly (re)formulated their prediction task as they discovered new information about data availability constraints and the predictability of alternative outcome variables. Based on their findings, Passi and Barocas (2019) describe problem formulation as a negotiated translation between high-level objectives and a concrete prediction task. Our work builds upon this finding to identify the more general evaluation desiderata and (re)formulation strategies driving the problem formulation process. We develop a process model through analysis of our findings and use this model to identify targeted opportunities to better support data scientists’ problem formulation practices going forward. For example, when viewed through the lens of our findings, the team studied by Passi and Barocas (2019) used one of five (re)formulation strategies (i.e., swapping) to change outcome variables when the first proved infeasible due to data availability constraints. We present design opportunities to help data scientists more effectively apply swapping ($\S$§ 4.2.3) and other (re)formulation strategies.
@@ 2.4. Toolkits Supporting Target Variable Construction
The limitations of data science toolkits and analysis methods have been documented across empirical studies examining data science practices (Jun et al., 2022a), software repositories (Hoess et al., 2025), and ML-based sciences (Kapoor and Narayanan, 2023). Jun et al. reviewed 20 commonly-used data analysis tools and found that the ecosystem of tools supports vastly different model implementations, even when using the same underlying mathematical equation (Jun et al., 2022a). Although these tools support nuanced model implementations, their low-level technical abstractions make it challenging for analysts to engage in the high-level conceptual reasoning required for hypothesis formalization (Jun et al., 2022a). As a result, interpretations of data, modeling assumptions, and implementation details can significantly influence outcomes and impact the validity of results (Hoess et al., 2025; Jun et al., 2022a, b; Kapoor and Narayanan, 2023). Research examining software engineering practices in code repositories, such as GitHub, further demonstrates how the choice of analysis tools can lead to substantial differences in simple metrics (e.g., commit counts and developer contributions) (Hoess et al., 2025; Harding, 2021; Lefever et al., 2021). This further poses threats to validity of findings derived from one analysis tool. Given this gap between technical abstractions and conceptual reasoning—and the embedded assumptions inherent in machine learning methods like feature engineering—it is not surprising that ML-based sciences continue to face a reproducibility crisis (Kapoor and Narayanan, 2023).
@@
To mitigate these issues, the HCI, ML, and FAccT research communities have developed a range of toolkits to intervene at points along data scientists’ inner and outer loop workflows. For instance, researchers have developed systems facilitating data collection and configuration (e.g., (Bhattacharya et al., 2024; Guo et al., 2011; Kandel et al., 2011; Rattenbury et al., 2017)), exploratory data analysis (e.g., (Wongsuphasawat et al., 2015; Kery et al., 2018)), model training (e.g., (Pang et al., 2020; Dingen et al., 2018)), caching intermediate states during data science experiments (e.g., (Schubotz et al., 2022)), and performance assessment (e.g., (Cabrera et al., 2023a; Yan et al., 2021)). Many tools have also been developed to help data scientists explain their models (e.g., (Ribeiro et al., 2018, 2016; Lundberg, 2017)) or assess fairness-related concerns (e.g., (Tramer et al., 2017; Bellamy et al., 2019; Weerts et al., 2023; Saleiro et al., 2018; Liu et al., 2024b; Cabrera et al., 2019; Ahn and Lin, 2019; Wexler et al., 2019; Dingen et al., 2018)). In the FAccT community, researchers have also developed responsible AI checklists and guidelines to help data scientists assess and document components of their AI system (e.g., dataset information (Gebru et al., 2021; Pushkarna et al., 2022), intended use cases (Mitchell et al., 2019)). To our knowledge, Gala et al. (2024) propose the only system specifically designed to support the target variable construction process. Instead, existing toolkits are often designed to support data scientists’ practices only after a target variable has been specified. This presents a pressing gap in data science support at a critical point in their workflows.
@@
However, prior HCI research suggests that such toolkits may prove ineffective if they are not designed around the existing practices of data scientists. For instance, past efforts to improve programming tools without a human-centered assessment of programmers’ needs yielded ineffective languages and systems (Myers et al., 2016; Hazzan and Tomayko, 2004; Seffah et al., 2005; Ko et al., 2011). Similarly, the utility of ML fairness toolkits was found to be limited in practice as they were designed for isolated use by technical roles, while actual fairness work often relied on deep cross-functional collaborations (Deng et al., 2022). Our work aims to help the research community avoid similar challenges while developing tools to support data scientists’ problem formulation processes. In particular, we identify opportunities for the HCI, CSCW, and ML research communities to develop tools that help data scientists more carefully weigh trade-offs across multiple evaluation criteria ($\S$§ 5.3.1), select ($\S$§ 5.3.2) and apply ($\S$§ 5.3.3) strategies, and evaluate the validity of their problem formulation more effectively ($\S$§ 5.3.4).

### v2

@@ Abstract.
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW447publicationmonth: 11doi: 10.1145/3757628ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learningccs: Computing methodologies Supervised learning
@@ 1. Introduction
Figure 1. An illustration of the relationship between a target variable, prediction task, and problem formulation. A prediction task describes a mapping from features to a target variable. A problem formulation connects a prediction task to high-level modeling goals, including factors such as how predictions are converted to decisions.
One critical aspect of problem formulation is the process of target variable construction (Figure 1). Many outcomes of interest to data scientists — such as the “authenticity” of student writing or the “healthcare need” of patients — are latent, theoretical constructs that are not directly observable. To develop a model targeting these constructs, data scientists must construct a proxy variable based on outcomes that are readily available. For example, a data scientist might choose to predict patients’ cost of medical care as a proxy for their “healthcare need” (Obermeyer et al., 2019). Traditionally, measurement theory assumes that analysts follow a top-down workflow while mapping unobserved constructs to measurements in data (Jacobs and Wallach, 2021). For example, a learning scientist might identify a latent construct they would like to measure among students (e.g., “math proficiency”) and design an operationalization (e.g., a test) before collecting data (e.g., student test scores). Disciplines in the quantitative social sciences have developed a rich set of guidelines (e.g., (Van’t Veer and Giner-Sorolla, 2016; Miller and Salkind, 2002)) and modeling tools (e.g., (Everett, 2013; Singh, 2007)) to help analysts implement this top-down measurement workflow.
@@
To understand data scientists’ current practices and challenges in constructing target variables—and to identify opportunities to better support target variable construction in practice—in this work, we conducted semi-structured interviews with fifteen data scientists from education (N=8) and healthcare (N=7). We asked data scientists to walk us through how they navigate target variable construction in prior modeling projects. Participants also engaged in a hypothetical target variable construction task, allowing us to observe their reasoning in a new modeling context. Our goal in these interviews was to understand how data scientists intertwine two conceptually distinct activities, measurement of latent constructs and prediction of target variables, during problem formulation (Mussgnug, 2022). While prior empirical studies of data science practice have often centered prediction (e.g., (Alspaugh et al., 2018; Wongsuphasawat et al., 2019; Kandel et al., 2012)), we broaden our scope of inquiry to investigate the interplay between measurement validity and predictive utility in problem formulation.
@@
Based on our findings, we synthesize a new process model that better reflects data scientists’ bricolage approach towards measurement in problem formulation (Figure 2). Data scientists balance the validity of their formulation with other criteria such as predictive performance, simplicity, portability, and resource requirements. Data scientists narrow in on a suitable formulation by drawing upon a rich repertoire of problem (re)formulation strategies. Data scientists apply a strategy, take a step back to assess their progress along each criterion, and then apply a new strategy until they believe all evaluation criteria are satisfied. Data scientists sometimes discontinue their projects if they conclude that no strategies are sufficient given the resources at hand.
@@
  • Empirical findings from interviews with fifteen data scientists in healthcare and education, synthesized into a process model that reveals how practitioners weigh competing criteria ($\S$ 4.1), apply five strategic approaches to (re)formulation ($\S$ 4.2), and evaluate validity through both theory-driven and data-driven practices ($\S$ 4.3).
  • Design implications for how the CSCW, HCI, and ML research communities might better scaffold data scientists’ bricolage practices going forward. These recommendations explore how we might enhance the validity of data scientists’ resulting formulations while also supporting the flexibility that makes bricolage effective ($\S$ 5).
@@ 2. Background and Related Work
In this section, we begin by introducing Claude Lévi-Strauss’ theory of Bricolage. We then describe existing Science and Technology Studies (STS), HCI, and CSCW research that has investigated questions related to categorization, measurement, and target variable construction in data science practice. We conclude by providing an overview of key concepts from measurement theory in the quantitative social sciences which we build upon in our interviews, findings, and discussion.
@@ 2.1. Bricolage and the Science of the Concrete
Bricolage, defined by Claude Lévi-Strauss as “making the most of available resources”(Levi-Strauss, 1966), describes how actors (i.e., bricoleurs) construct artifacts in creative and adaptive ways. Bricolage is central to what Lévi-Strauss calls the science of the concrete, which contrasts with modern scientific thought by focusing on the materials directly at hand rather than abstract principles. Bricolage contains three elements: repertoire, dialogue, outcome (Duymedjian and Rüling, 2010). Repertoire describes the resources—physical artifacts, ideas, or skills—that the bricoleur brings to bear on a problem. Whereas the engineer or scientist acquires resources tailor-made to specification, the bricoleur assembles “odds and ends” that might be useful in the future. Then, when faced with a new problem, the bricoleur engages in dialogue with these materials—a process of arranging and re-arranging materials in her repertoire to find a workable solution. While engaging in dialogue, the bricoleur adopts old objects for new purposes and combines concepts with materials in novel ways. Finally, outcome describes the artifact resulting from the bricolage process. While bricolage entails a tolerance for imperfection, bricoleurs sometimes identify “brilliant unforeseen results” (Levi-Strauss, 1966) as part of their process.
Prior research has used Lévi-Strauss’ framework to describe many forms of knowledge work as bricolage (e.g., entrepreneurship (Mateus and Sarkar, 2024; Duymedjian and Rüling, 2010), design (Büscher et al., 2001; Louridas, 1999; Vallgårda and Fernaeus, 2015), management (Duymedjian and Rüling, 2010), and programming (Turkle, 2011)). For instance, Baker and Nelson (2005) use bricolage to explain how entrepreneurs build business ventures under resource constraints. They observe that entrepreneurial success hinges on the ability of firms to combine their limited existing resources in creative, adaptive ways.
@@
Turkle describes this style of tinkering as important for developing an intuition for the complexity of computer programs. Like a painter who steps back to inspect a canvas between brush strokes, programmers try one approach, re-evaluate, and then try another. This process often leads programmers to more elegant solutions than could be imagined through structured planning.
In this work, we draw a novel connection between data science work and bricolage. Much like an entrepreneur, programmer, or painter, data scientists construct target variables by combining existing resources in creative, adaptive ways. Instead of instantiating abstract plans through top-down design—akin to the structured programming approach taught in classrooms—data scientists engage in more direct dialogue with their data and modeling tools, making do with the materials at hand. And analogous to past research investigating how to support such exploratory and improvisational coding practices (Kery and Myers, 2017), this has implications for the design of practical tools to support target variable construction in practice. We highlight connections between target variable construction and bricolage throughout the findings ($\S$ 4) and describe implications for tool design in the discussion ($\S$ 5).
@@ 2.2. Tracing the Social Construction of Target Variables
Target variable construction also relates to a rich body of Science and Technology Studies (STS) and philosophy of science literature that investigates the social processes that give rise to categories in data (Latour and Woolgar, 2013; Bowker, 2000; Pine and Liboiron, 2015; Harvey et al., 2024; Porter, 1996; Crosby, 1997; Hacking, 1999; Mol, 2002; Suchman, 1993). Latour and Woolgar (2013) describe the social construction of scientific facts within laboratories, highlighting that facts are not merely discovered but actively produced through a complex interplay of social and technical factors. Likewise, Bowker (2000) examines how the International Classification of Diseases (ICD), shape and are shaped by the social and political contexts in which they are embedded. Building on Bowker’s work, recent studies have also explored the construction of racial and ethnic categories in AI training data (e.g., (Abdu et al., 2023; Mickel, 2024)) and charted measurement assumptions underpinning motion capture technologies (Harvey et al., 2024).
Most related to our study, Muller et al. (2021) interviewed fifteen data scientists to investigate the labeling practices of data science teams. Muller et al. (2021) identified three approaches that teams used while defining labels: principled, iterative, and improvisational. Teams engaging in principled design followed a top-down planning approach, systematically specifying annotation guidelines before data labeling. Teams leveraging iterative design took a bottom-up approach, incorporating opportunities for annotation protocol refinement throughout the labeling process. Teams engaging in improvisational design viewed data labeling as an open-ended, exploitative process. Our study investigates a distinct question of how data science teams construct target variables for predictive modeling tasks. Rather than studying how data science teams design “ground truth” labels anew, we examine how data scientists connect their existing data to their high-level modeling objectives.
@@ 2.3. Empirical Studies of Data Science Practice
However, we currently have an incomplete understanding of how data scientists construct target variables while formulating prediction tasks. To date, Passi and Barocas (2019) provide the most detailed account of how problem formulation unfolds in real-world modeling contexts. Over the course of an eighteen-month ethnographic study of a corporate data science team, Passi and Barocas (2019) observed that data scientists repeatedly (re)formulated their prediction task as they discovered new information about data availability constraints and the predictability of alternative outcome variables. Based on their findings, Passi and Barocas (2019) describe problem formulation as a negotiated translation between high-level objectives and a concrete prediction task. Our work builds upon this finding to identify the more general evaluation desiderata and (re)formulation strategies driving the problem formulation process. We develop a process model through analysis of our findings and use this model to identify targeted opportunities to better support data scientists’ problem formulation practices going forward. For example, when viewed through the lens of our findings, the team studied by Passi and Barocas (2019) used one of five (re)formulation strategies (i.e., swapping) to change outcome variables when the first proved infeasible due to data availability constraints. We present design opportunities to help data scientists more effectively apply swapping ($\S$ 4.2.3) and other (re)formulation strategies.
@@ 2.4. Toolkits Supporting Target Variable Construction
The limitations of data science toolkits and analysis methods have been documented across empirical studies examining data science practices (Jun et al., 2022a), software repositories (Hoess et al., 2025), and ML-based sciences (Kapoor and Narayanan, 2023). Jun et al. reviewed 20 commonly-used data analysis tools and found that the ecosystem of tools supports vastly different model implementations, even when using the same underlying mathematical equation (Jun et al., 2022a). Although these tools support nuanced model implementations, their low-level technical abstractions make it challenging for analysts to engage in the high-level conceptual reasoning required for hypothesis formalization (Jun et al., 2022a). As a result, modeling assumptions and interpretations of data can impact the validity of results (Hoess et al., 2025; Jun et al., 2022a, b; Kapoor and Narayanan, 2023). Research examining software engineering practices in code repositories, such as GitHub, further demonstrates how the choice of analysis tools can lead to substantial differences in simple metrics (e.g., commit counts and developer contributions) (Hoess et al., 2025; Harding, 2021; Lefever et al., 2021). This further poses threats to validity of findings derived from one analysis tool. Given this gap between technical abstractions and conceptual reasoning—and the embedded assumptions inherent in machine learning methods like feature engineering—it is not surprising that ML-based sciences continue to face a reproducibility crisis (Kapoor and Narayanan, 2023).
@@
To mitigate these issues, the HCI, ML, and FAccT research communities have developed a range of toolkits to intervene in data scientists’ inner and outer loop workflows. For instance, researchers have developed systems facilitating data collection and configuration (e.g., (Bhattacharya et al., 2024; Guo et al., 2011; Kandel et al., 2011; Rattenbury et al., 2017)), exploratory data analysis (e.g., (Wongsuphasawat et al., 2015; Kery et al., 2018)), model training (e.g., (Pang et al., 2020; Dingen et al., 2018)), caching intermediate states during data science experiments (e.g., (Schubotz et al., 2022)), and performance assessment (e.g., (Cabrera et al., 2023a; Yan et al., 2021)). Many tools have also been developed to help data scientists explain their models (e.g., (Ribeiro et al., 2018, 2016; Lundberg, 2017)) or assess fairness-related concerns (e.g., (Tramer et al., 2017; Bellamy et al., 2019; Weerts et al., 2023; Saleiro et al., 2018; Liu et al., 2024b; Cabrera et al., 2019; Ahn and Lin, 2019; Wexler et al., 2019; Dingen et al., 2018)). In the FAccT community, researchers have also developed responsible AI checklists and guidelines to help data scientists assess and document components of their AI system (e.g., dataset information (Gebru et al., 2021; Pushkarna et al., 2022), intended use cases (Mitchell et al., 2019)). To our knowledge, Gala et al. (2024) propose the only system specifically designed to support the target variable construction process. Instead, existing toolkits are often designed to support data scientists’ practices only after a target variable has been specified. This presents a pressing gap in data science support at a critical point in their workflows.
@@
However, prior human–computer interaction research suggests that such toolkits may prove ineffective if they are not designed around the existing practices of data scientists. For instance, past efforts to improve programming tools without a human-centered assessment of programmers’ needs yielded ineffective languages and systems (Myers et al., 2016; Hazzan and Tomayko, 2004; Seffah et al., 2005; Ko et al., 2011). Similarly, the utility of ML fairness toolkits was found to be limited in practice as they were designed for isolated use by technical roles, while actual fairness work often relied on deep cross-functional collaborations (Deng et al., 2022). Our work aims to help the research community avoid similar challenges while developing tools to support data scientists’ problem formulation processes. In particular, we identify opportunities for the HCI, CSCW, and ML research communities to develop tools that help data scientists more carefully weigh trade-offs across multiple evaluation criteria ($\S$ 5.3.1), select ($\S$ 5.3.2) and apply ($\S$ 5.3.3) strategies, and evaluate the validity of their problem formulation more effectively ($\S$ 5.3.4).